a) A given electrical circuit operating on "V" volts with "R" ohms of resistance draws 5.8 amps of current. How much current will it draw if "V" is multiplied by 2.7?

b) How much current will the circuit above draw if it is connected to the original "V" but the resistance is multiplied by 5.0?
Answer #1

Part A.

Using Ohm's law:

E = I*R

When E = V & R = R, then

I = E/R = V/R = 5.8 Amp

Now when E1 = 2.7*V, then

I1 = E1/R = 2.7*V/R

I1 = 2.7*I = 2.7*5.8 = 15.66 A

Part B.

Now when R1 = 5.0*R

then

I1 = E/R1 = V/(5.0*R) = (1/5.0)*(V/R)

I1 = (1/5.0)*5.8 = 5.8/5.0 = 1.16 Amp

Please Upvote.